OpenStudy (anonymous):

The sum of a number and 6 times its reciprocal is 5. Find the number. If the problem has two solutions, write down your answers in the box below

OpenStudy (anonymous):

6*reciprical = 5?

OpenStudy (anonymous):

6/5?

OpenStudy (anonymous):

Solve the following for x:\[(x+6)\left(\frac{1}{x}\right)=5\]\[x=\frac{3}{2} \]